# A Geometric Tension Dynamics Model of Epithelial Convergent Extension

### v2

@@ Results
With the LTC order parameter and the T1 threshold in hand, we can now quantify the dynamics of tensions in the simulations (see Fig. 5A) and experiments (see companion paper Ref. [17] and Fig. S2). Because isogonal strain shifts the T1 threshold, it will have a significant effect on the LTC order parameter distribution. We therefore performed simulations where we impose the isogonal strain observed in the Drosophila germ band [17] (see SI Sec. 2 for details). The initial tension triangulations where taken to be perturbed lattices with tension anisotropy $0.175$. The germ band is stretched by the invagination of the adjacent mesoderm tissue causing isogonal strain along the axis of tension anisotropy (dorsal–ventral). This shifts the T1 threshold towards stronger tension anisotropy $\tilde{s}$ (cf. Fig. 4E).
At the beginning of the simulation, the local tension anisotropy $\tilde{s}$ is small and there is no bias towards cables or bridges. Positive tension feedback amplifies local anisotropy and thus drives the tension configurations towards the T1 threshold (Fig. 5A). At the onset of T1s, we observe an increased fraction of tension bridges, in agreement with the experimental observations (8 min in Fig. 5A). At late times, where the tissue becomes highly disordered, the distribution in LTC space shifts more towards tension cables (30 min). As we will see below, this late time distribution is reproduced by a random Delaunay triangulation. Time traces of the median anisotropy $\tilde{s}$ and (weighted) median LTC phase $\tilde{\psi}$ show qualitative agreement with data from the Drosophila germ band [17], see Fig. 5B. In fact, a quantitative agreement can be achieved by shifting the median anisotropy and LTC phase by constant offsets, as indicated by the dashed lines in Fig. 5B. These offsets may be a consequence of noise in the experimental data, which we support in Fig. S8 with simulations that incorporate Langevin noise in the tension dynamics.
To understand how tension bridges emerge transiently, consider the shape dynamics of a single, isolated tension triangle governed by winner-takes-all feedback Eq. (5). Starting from a configuration with nearly equal tensions, i.e. a nearly equilateral tension triangle, winner-takes-all feedback causes the highest tension (longest edge in the triangle) to grow at the expense of the other two. The triangle is thus driven towards an increasingly obtuse shape, as illustrated in Fig. 5C. Visualizing this dynamics as a flow in LTC space, shows that it drives the tension configurations towards the T1 threshold and thereby causes the cell rearrangements. The single-triangle simulation successfully predicts early dynamics of the LTC distribution until the onset of cell rearrangements (see Fig. 5B, C’). While positive tension feedback explains the emergence of tension bridges at the single-vertex level, it is not enough to produce an alternating pattern of tensions across cells. This pattern requires that the elementary tension configuration motifs (bridges) fit together coherently, i.e. their tension anisotropy is aligned across cells. This requires that the coordination number of a majority of cells is 6, i.e. that most cells are hexagons. This explains why hexagonal packing order is required to drive coherent T1s that underlie rapid convergent extension.
For a highly ordered initial packing, i.e. a nearly hexagonal lattice of cells, the lattice orientation relative to the axis of mean tension anisotropy determines the initial fraction of tension bridges. It is maximal when one side of the hexagonal cells is parallel to the orientation of tension anisotropy and minimal when one side is perpendicular (see Fig. S7). In simulations we find that this initial bias only weakly affects the initial rate of tissue extension and has no significant effect at later times.
In the simulation discussed above we imposed isogonal strain to account for the transient stretching of the germ band by the invaginating mesoderm. In simulations without imposed isogonal strain, the bridge fraction does not transiently increase (see Fig. S8). This is because the T1 threshold is at a significantly lower anisotropy for tension bridges than for tension cables and this difference is more pronounced for vanishing isogonal strain, causing tension bridges to be rapidly eliminated by T1s (see gray line in Fig. 5C). We therefore predict that T1s will happen at a lower tension anisotropy (and lower tension-bridge fraction) in a tissue with no isogonal strain, e.g. in a twist mutant embryo where mesoderm invagination is abolished.
Loss of order in local tension configurations.
@@
Above, we have seen that the cell array becomes disordered as cells rearrange. The coordination number statistics and average shape index approach a random Voronoi tessellation (cf. Fig. 2E, F). This suggests that we can characterize the corresponding tension triangulations as random Delaunay triangulations. Their statistics of such a triangulation depend on the underlying stochastic point process. We use the same hard disk sampling method as above, controlled by the packing fraction $\rho$ to generate a family of Delaunay triangulations. The triangle shape statistics (characterized by the LTC parameter) found at late times both in simulations and in the Drosophila germ band are reasonably well reproduced by a random Delaunay triangulation seeded from a hard disks with $\rho\approx 0.2$ (Fig. 5D, D’ and Fig. S3). This observed “randomization” in tension space suggests that the triangle edge flips can be statistically understood as a random “mixing”. Notably, the late-time distribution exhibits a bias towards tension cables. The loss of tension bridges causes active T1s to become incoherent and incompatible between adjacent cells, contributing to a slowdown of tissue extension found in tissue scale simulations and in the germ band [17].
@@
Taken together, we find that the time course of LTC distribution agrees between the model and the experimental data. Next, we show how changing various aspects of the model affects the LTC distribution, highlighting that the LTC parameter can be used to distinguish different tension dynamics based on statistical signatures of cell-scale observations.
@@
Figure 6: Saturating tension feedback causes tension cable formation and, hence, fails to drive convergent extension. A Local tension configurations are driven towards tension cables due to saturation of positive feedback. Compared to winner-takes-all feedback, this resulting tension dynamics is significantly less efficient at the local configuration towards the T1 threshold (cf. Fig. 5C). B In a tissue simulation, saturating feedback generates only little convergent extension (quantification in D). The initial configuration is shown semi-transparent in the background. Cell color indicates coordination number (cf. Fig. 2A) C The late-time LTC distribution shows a strong cable bias and differs significantly from the random Delaunay distribution emerging in simulations with winner-takes-all feedback (cf. Fig. 5D. D–F Time series quantification showing that saturating feedback (teal), compared to winner-takes-all feedback (purple), yields very little aspect ratio change (D), a significantly reduced T1 rate (E) and a strongly increased fraction of tension cables (F). Note that in contrast to Fig. 5, the simulations were performed without an imposed isogonal strain such that no significant formation of tension bridges is observed for winner-takes-all feedback.
So far, we used a “winner-takes-all” local tension feedback mechanism, Eq. (5), which drives the formation of tension bridges as illustrated in Fig. 5C. In contrast, when positive feedback rapidly saturates, adjacent high tension interfaces no longer compete, thus leading to the formation of tension cables (Fig. 6A; see SI Sec. 2 for details). The trajectories in LTC space obtained from single-triangle simulations show that saturating feedback is less efficient at driving the local tension configuration towards the T1 threshold. Indeed, tissue scale simulations with such feedback produce very little convergent extension (see Fig. 6B, D and Movie 3). The rate of T1 transitions is significantly reduced (Fig. 6E), and in contrast to “winner-takes-all” feedback, a significant fraction of T1 transitions (approximately 20%) are reversible, i.e. the newly formed edge rapidly re-collapses (see SI Sec. 2). Saturating tension feedback might therefore explain the reversible T1s observed in certain Drosophila mutants [38]. Furthermore, the LTC distribution develops a significant bias towards tension cables as shown in Fig. 6C and Fig. 6F. Such persistent tension cables are observed in Drosophila abl mutants, suggesting that knockout of abl might impair positive tension feedback (see Discussion). Taken together, the above findings show that flow in LTC parameter space obtained from single-triangle simulations serves as a simple tool to predict the cell scale behavior (efficiency of active T1s, emergence of tension cables vs bridges) for a given tension-feedback law.
@@
Figure 7: Combining active and passive tissue regions. A, B The ellipsoidal geometry of the Drosophila embryo (A) is mimicked by a simplified simulation geometry corresponding to an unrolled cylinder (B), whose azimuthal axis corresponds to the dorso-ventral (vertical) axis of the embryo. The different behaviour of the dorsal amnioserosa (AS) and the lateral ectoderm (LE) tissue is represented by the passive and active region in the simulation domain respectively. C Positive feedback in the active region amplifies an initial DV anisotropy of tension and thus drives extension along the AP axis. Since the embryos circumference is fixed (implemented via a slip wall at the dorsal boundary), the passive region is stretched along the DV axis. Only half of the simulation domain is shown, corresponding to one lateral side of the left-right symmetric embryo. Three-by-three patches of cells are highlighted to show cell rearrangements (cf. Movie 4). D Tissue deformation, as measured by the change in aspect ratio of the active region, $a(t)/a(0)$, stalls after an elongation (and perpendicular contraction) by a factor $\sim 1.6$. Shading indicates standard deviation over $N=3$ simulation runs of $N_{\mathrm{cells}}\approx 750$ cells. E Trajectories of cell centroids showing the tissue scale flow, resembling the characteristic flow of Drosophila germ-band extension [2].
So far we have considered a patch of active cells with free boundary conditions. However, the epithelium of the early Drosphila embryo forms a closed surface that is approximately an elongated ellipsoid. Therefore, deformation of one tissue region has to be compensated by an opposite deformation of another region. For example, on the Drosphila embryo, the dorsal amnioserosa is passively stretched along the dorso-ventral (DV) axis and compressed along the anterior-posterior (AP) axis to compensate the convergence extension of the germ band. In the following, we investigate the interplay of active and passive tissue deformations.
@@
To mimic the cylindrical geometry of the embryo’s trunk (Fig. 7A) we simulated a rectangular tissue patch with “slip walls” at the top and bottom boundary (Fig. 7B). Along the slip walls, cell centroids, marked by black disks, are restricted to move along the wall. These boundary conditions fix the DV extent of the tissue, corresponding to the fixed circumference of the embryo. We divide the tissue into active and passive regions to account for the different mechanical properties of the lateral ectoderm and the dorsal tissue which becomes the amnioserosa [2, 17]. (The simulation domain is symmetric with respect to reflection across the $x$-axis, corresponding to the left-right symmetry of the embryo.) Cortical tensions are governed by positive feedback in the active region and by tension homeostasis in the passive region. Further, passive cells (subscript $p$) are taken to be soft $\mu_{p}=0.2\mu_{a},\;\lambda_{p}=0.2\lambda_{a}$ compared to active cells [39]. In addition, we allow interface angles in the passive region to slightly deviate from those imposed by the tension triangulation, reflecting the fact that the overall scale of cortical tensions is lower in the passive tissue [2]. We initialize the simulation with a slightly perturbed hexagonal packing of cells (motivated by the experimental observations, see Fig. S1) and the experimentally observed tension anisotropy aligned along the DV axis [17].
@@
Starting from this initial condition, the simulation reproduces salient features of the tissue-scale dynamics in the embryo (see Fig. 7C and Movie 4). In the active region (“lateral ectoderm”) active cell rearrangements drive tissue extension along the AP axis and contraction along the DV axis. The passive region (“amnioserosa”) is stretched along the DV axis, accommodating the fixed circumference of the embryo. Notably, this stretching leads to T1s in the passive region as is visible from the highlighted cells in Fig. 7C. On the tissue level, the coupling of active and passive regions gives rise to the tissue flow pattern characteristic of Drosophila germ-band elongation [2] as shown in Fig. 7E.
@@
Figure 8: Tissue flow requires cell rigidity and large-scale genetic patterning. A Cartoon of an an embryo without dorso-ventral mechanical patterning, such as a Toll[RM9] mutant where cells all around the circumference adopt a lateral ectoderm (LE) fate. A’ Simulation of a tissue patch without a passive region. While active T1s deform the tension triangulation (top), the cell array (bottom) is blocked from elongating by the fixed DV “circumference” (implemented via slip walls). Instead, cell rearrangements are compensated by isogonal cell elongation (quantified in E), leading to a build-up of elastic energy. B The length dynamics of collapsing and emerging interfaces do not significantly depend on whether the tissue as a whole deforms (WT) or is blocked from undergoing convergent extension (no DV modulation). Purple band shows one standard deviation. Time 0 is the moment of interface collapse. C Orientation of collapsing and emerging interfaces showing that only the orientation of the latter is qualitatively affected by tissue-scale modulation of activity. The orientation of emerging interfaces is slightly biased along the DV direction in the absence of a passive tissue, while it is biased in AP direction in the WT case. D Convergent extension is strongly suppressed in absence of DV modulation of activity (“No DV”) and when active cells have vanishing shear modulus ($\mu_{a}=0$; see F and G). $a(t)$ measures the aspect ratio of the active tissue region. E Histograms of cell shape elongation measured by relative difference, $(S_{\text{DV,DV}}-S_{\text{AP,AP}})/(S_{\text{DV,DV}}+S_{\text{AP,AP}})$, of the AP-AP and DV-DV components of the shape tensor $S$. In the WT case (cf. Fig. 7), cells remain nearly isotropic while they become significantly oriented along the DV axis when DV modulation of activity or rigidity of active cells are abolished. F When the active cells have no shear rigidity ($\mu_{a}=0$), cell rearrangements are compensated by isogonal cell elongation (quantified in E) without incurring an elastic energy build-up. Thus, almost no tissue scale convergent-extension takes place. G Net amount of convergent extension as a function of the shear modulus of active cells shows that active cells need to be stiffer than the surrounding passive tissue for active T1s to drive efficient convergent extension.
@@
The total tissue extension found in the simulations that combine active and passive tissue regions is smaller than the extension of active tissue patches with free boundaries (compare Fig. 7D and Fig. 2C). This suggests that the passive tissue resists deformation. Indeed, cells in the active region are slightly elongated along the DV axis, indicating that the passive tissue pulls on them. In the following, we further investigate the role of the spatial modulation of the cells’ mechanical properties along the DV axis. Figure 8A’ shows a simulation without DV modulation where all cells are active. Positive tension feedback drives active T1s everywhere, as is manifest in the deformation of the tension triangulation (Fig. 8A’, right). However, because of the slip-wall boundary conditions, the tissue cannot contract along the DV axis so that T1s do not result in convergence extension (see Movie 5; quantification in Fig. 8D). Instead, cell rearrangements are compensated by isogonal deformations resulting in elongated cell shapes (as quantified in Fig. 8E). We predict that this scenario will be realized in Toll[RM9] mutant embryos where all cells around the embryo’s circumference adopt a ventro-lateral fate [33], as illustrated in the cartoon in Fig. 8A.
The stretching of cells leads to a buildup of elastic energy (see Fig. S10) which eventually leads to convergence issues in the numerical simulation for $t\gtrsim 20\;\mathrm{min}$. In Toll[RM9] mutant embryos, some of this elastic energy is released by the formation of folds (buckling) [33] which our 2D simulations cannot capture. A partial compensation of cell rearrangements by cell shape changes might be observed in the germ band of mutants where the soft amnioserosa is abolished (e.g. in dpp[hin46] mutants).
@@
An interesting observation from the above simulations is that the length dynamics of collapsing and emerging interfaces is not significantly affected by the lack of tissue-scale mechanical patterning (Fig. 8B), even though there is no tissue extension. Interface elongation in the absence of tissue elongation has previously been observed in experiments where germ band extension has been blocked by cauterization near the posterior pole [8, 40]. We mimic the experiments of Ref. [8] by adding slip walls to the anterior and posterior boundary of the simulated tissue (Fig. S9) to block tissue extension. In this scenario, we also observe that newly formed interfaces extend, but with a slightly different dynamics than in WT. In further agreement with experimental observations, we find that the orientation of extending interfaces is no longer biased along the AP axis when tissue extension is blocked (see Fig. 8C). Since interface extension in our model is a purely passive process, we conclude that no active mechanisms – such as medial myosin pulses [8] – are necessary for the elongation of new interfaces. Instead, interface elongation results from the fundamental temporal asymmetry of the intercalation process, i.e. the low level of active tension on the new interface (see companion paper [17]).
@@
Rather than abolishing the DV modulation of activity, one can also change the passive elastic properties of the cells. Recall that in the model, the resistance of cells against deformations is described by the cell-shape elastic energy, Eq. (3) parameterized by the Lamé coefficients $\lambda$ (resistance to area changes) and $\mu$ (resistance to shear deformation). Fig. 8F shows a simulation where the shear modulus $\mu_{a}$ of active cells is set to zero, so that the cells do not resist (area-preserving) elongation. Active T1s can therefore be fully compensated by cell elongation through isogonal deformations without incurring an elastic energy. As a result, there is no net tissue deformation (see Movie 5). In other words, cell-shape rigidity is required to maintain rotund cell shapes (i.e. resist isogonal shear deformations) and thus translate active T1s into net tissue deformation. The tissue deformation by isogonal modes is determined by a balance of external forces and internal resistance of cells the shape changes. Here, the external forces acting on the active tissue result from the passive tissue’s resistance to deformation, which is in turn set by shear modulus $\mu_{\mathrm{p}}$. The ratio of the shear moduli in the active vs the passive region, $\mu_{\mathrm{a}}/\mu_{\mathrm{p}}$, determines how much the active region deforms (see Fig. 8G). Only when the cells in the active tissue are more rigid than those in the passive region ($\mu_{\mathrm{a}}/\mu_{\mathrm{p}}>1$), is it energetically favorable to isogonally deform the passive region rather than the active region. This predicts that GBE can be impaired by stiffening the dorsal tissue (amnioserosa).

### v3

@@ Results
The LTC order parameter and the T1 threshold in hand, we can quantify the dynamics of tensions in the simulations (see Fig. 4A) and experiments (see companion paper Ref. [Brauns.etal2024elife] and Fig. S2). Because isogonal strain shifts the T1 threshold (cf. Fig. 3E), it will have a significant effect on the LTC order parameter distribution. We imposed in our simulations the isogonal strain observed in the Drosophila germ band [Brauns.etal2024elife], where invagination of the adjacent mesoderm tissue causes isogonal strain along the axis of tension anisotropy (see SI Sec. 2).
LTC histograms show an increase in anisotropy and a transient bias towards tension bridges, before convergence to a steady state biased towards tension cables as the tissue becomes disordered (Fig. 4A). Time traces of the median anisotropy $\tilde{s}$ and (weighted) median LTC phase $\tilde{\psi}$ show qualitative agreement with the experimental data from the Drosophila germ band [Brauns.etal2024elife] (Fig. 4B). Quantitative agreement can be achieved by adding constant offsets (dashed lies in Fig. 4B), which may be a consequence of noise in the experimental data (see SI-Fig. S9 for simulations incorporating Langevin noise).
Figure 4: Dynamics of LTC order in simulations and the Drosophila germ band. A Heat maps showing the distribution of local tension configurations in a simulation with imposed isogonal strain matching the experimental observations. Data aggregated from $N=6$ simulation runs of $\sim 10^{3}$ cells each. B Driven by winner-takes-all feedback, the magnitude of tension anisotropy magnitude $\tilde{s}$ (top) and the tension bridge fraction (measured by the LTC phase $\tilde{\psi}$; bottom) increase. As the tissue becomes disordered due to cell rearrangements, the bridge fraction starts decreasing at ca. 10 min. Solid lines show the median of the LTC distributions where the phase $\tilde{\psi}$ is weighted with the magnitude $\tilde{s}$. The width of bands showing the standard error is comparable to the line width. Shifting the median from simulations by a constant offset (dashed orange lines) yields a quantitative match to the experimental data. C Shape dynamics of a single tension triangle driven by winner-takes-all feedback rapidly drives the tensions towards the T1 threshold with a slight bias towards tension bridges.
To understand the LTC dynamics, consider the shape dynamics of a single, isolated tension triangle governed by winner-takes-all feedback, Eq. 5]. Starting from a configuration with nearly equal tensions, the highest tension grows at the expense of the other two, driving the triangle towards an increasingly anisotropic and obtuse shape, as illustrated in Fig. 4C. This LTC flow drives the tension configurations towards the T1 threshold and thereby causes cell rearrangements. The single-triangle simulation successfully predicts early dynamics of the LTC distribution until the onset of cell rearrangements (Fig. 4B). The single-triangle picture also highlights the impact of isogonal strain; in simulations without imposed isogonal stretching, the T1–threshold is positioned so that bridges are rapidly eliminated by T1s (gray line in Fig. 4C), and no transient bride bias is observed. We predict that this will occur in twist or snail mutant embryos where mesoderm invagination is abolished.
While positive tension feedback explains the emergence of tension bridges at the local (single-triangle) level, it is not enough to produce an alternating pattern of tensions across cells. For such a pattern, tension bridges must fit together coherently, i.e. their tension anisotropy is aligned across cells. This requires that the coordination number of a majority of cells is 6, i.e. that most cells are hexagons. This explains why some degree of hexagonal packing order is required to drive coherent T1s that underlie rapid CE. Notably, however, long-range hexatic order is not necessary, as the local positive feedback on tensions is able to promote the local alignment of hexatic and nematic order in tension space, as manifest in the LTC phase dynamics.
@@
In Fig. 2, we have seen that as cells rearrange, the cell array becomes disordered with the coordination number statistics approaching a random Voronoi tessellation. This suggests that the corresponding tension triangulations resemble random Delaunay triangulations. To generate a family of such random triangulations, we use the same hard disk sampling method as above, controlled by the packing fraction $\rho$. We find that the triangle shape (LTC) statistics of a random Delaunay triangulation with $\rho\approx 0.2$ reproduces the late time statistics observed in simulations and in the Drosophila germ (see Fig. S3). Notably, the late-time distribution exhibits a slight bias towards tension cables. The loss of tension bridges causes active T1s to become incompatible between adjacent cells, contributing to a slowdown of tissue extension found in tissue scale simulations and in the germ band [Brauns.etal2024elife].
@@
Taken together, we find that the time course of LTC distribution agrees between the model and the experimental data. Next, we show how changing aspects of the model affects the LTC distribution, highlighting that the LTC parameter can be used to distinguish different tension dynamics based on statistical signatures of cell-scale observations.
@@
The “winner-takes-all” local tension feedback mechanism, Eq. [5], considered so far is efficient at driving T1s because it causes the formation of tension bridges as illustrated in Fig. 4C. In contrast, when positive feedback rapidly saturates, adjacent high tension interfaces no longer compete, leading to the formation of tension cables (Fig. 5A; see SI Sec. 2 for details). The trajectories in LTC space obtained from single-triangle simulations show that saturating feedback is less efficient at driving the local tension configuration towards the T1 threshold. Indeed, tissue scale simulations with such feedback produce very little CE (see Figs. 5B and 5D and Movie 3). The rate of T1 transitions is significantly reduced (Fig. 5D), and in contrast to “winner-takes-all” feedback, a significant fraction of T1 transitions (approximately 20%) are reversible, i.e. the newly formed edge rapidly re-collapses (see SI Sec. 2). Indeed, the T1 rate is transiently quite high but there is very little CE, suggesting that T1s along cables are inefficient at driving tissue deformation. Saturating tension feedback might therefore explain the reversible T1s observed in certain Drosophila mutants [Bardet.etal2013]. As predicted from single-triangle shape-space flow (Fig. 5A) the LTC distribution develops a significant bias towards tension cables as shown in Fig. 5C. Such persistent tension cables are observed in Drosophila abl mutants, suggesting that knockout of abl might impair positive tension feedback (see Discussion).
The above findings show that the LTC order parameters capture important structural features on the cellular scale that strongly affect the dynamics and efficiency of T1s processes. Flow in LTC parameter space obtained from single-triangle simulations serves as a simple tool to predict the cell scale behavior (efficiency of active T1s, emergence of tension cables vs bridges) for a given tension-feedback law.
Figure 5: Saturating tension feedback causes tension cable formation and, hence, fails to drive CE. A The flow in local tension configuration space induced by saturating positive feedback is inefficient at driving T1s. B Saturating feedback generates only little CE (quantification in D). The initial configuration is shown semi-transparent in the background. Cell color indicates coordination number (cf. Fig. 2A). C The late-time LTC distribution shows a strong cable bias and differs significantly from the random Delaunay distribution emerging in simulations with winner-takes-all feedback (cf. Fig. 4D). D Saturating feedback (teal), compared to winner-takes-all feedback (purple), yields very little aspect ratio change and a significantly reduced T1 rate.
@@
The epithelium of the early Drosphila embryo forms a closed, approximately ellipsoidal surface. Therefore, the deformation of one tissue region has to be compensated by an opposite deformation elsewhere. Specifically, the dorsal amnioserosa is passively stretched along the dorso-ventral (DV) axis and compressed along the anterior-posterior (AP) axis to compensate the convergence extension of the germ band. To investigate this interplay of active and passive tissue deformations, we mimic the cylindrical geometry of the embryo’s trunk (Fig. 6A) by a rectangular tissue patch with “slip walls” at the top and bottom boundary (Fig. 6B). Along the slip walls cell centroids are restricted to move along the wall, thus fixing the DV extent (i.e. “circumference”) of the tissue. To account for the different mechanical properties of the lateral ectoderm and the dorsal tissue, we divide the tissue into active and passive regions [Streichan.etal2018, Brauns.etal2024elife]. In the former, cortical tensions are governed by positive feedback in the active region while tension homeostasis is imposed in the latter. Further, passive cells (subscript $p$) are taken to be soft $\mu_{p}=0.2\mu_{a},\;\lambda_{p}=0.2\lambda_{a}$ compared to active cells [Rauzi.etal2015]. In addition, we allow interface angles in the passive region to slightly deviate from those imposed by the tension triangulation, reflecting the fact that the overall scale of cortical tensions is lower in the passive tissue [Streichan.etal2018]. We initialize the simulation with a slightly perturbed hexagonal packing of cells and the experimentally observed tension anisotropy aligned along the DV axis [Brauns.etal2024elife].
@@
Starting from this initial condition, the simulation reproduces salient features of the tissue-scale dynamics in the embryo (see Fig. 6C and Movie 4). In the active region (“lateral ectoderm”, LE) active cell rearrangements drive tissue extension along the AP axis and contraction along the DV axis. The passive region (“amnioserosa”, AS) is stretched along the DV axis, accommodating the fixed circumference of the embryo. Notably, this stretching leads to T1s in the passive region as is visible from the highlighted cells in Fig. 6C. On the tissue level, the coupling of active and passive regions gives rise to the tissue flow pattern characteristic of Drosophila germ-band elongation [Streichan.etal2018] as shown in Fig. 6D.
Figure 6: Combining active and passive tissue regions. A, B The ellipsoidal geometry of the Drosophila embryo (A) is mimicked by a simplified simulation geometry corresponding to an unrolled cylinder (B), whose azimuthal axis corresponds to the dorso-ventral (vertical) axis of the embryo. The different behavior of the dorsal amnioserosa (AS) and the lateral ectoderm (LE) tissue is represented by the passive and active regions in the simulation domain respectively. C Positive feedback in the active region amplifies an initial DV anisotropy of tension and thus drives extension along the AP axis. Since the embryo’s circumference is fixed (implemented via a slip wall at the dorsal boundary), the passive region is stretched along the DV axis. Only half of the simulation domain is shown, corresponding to one lateral side of the left-right symmetric embryo. Three-by-three patches of cells are highlighted to show cell rearrangements (cf. Movie 4). D Trajectories of cell centroids showing the tissue scale flow, resembling the characteristic flow of Drosophila germ-band extension [Streichan.etal2018].
@@
Figure 7: Tissue flow requires cell rigidity and large-scale genetic patterning. A Cartoon of an an embryo without dorso-ventral mechanical patterning, such as a Toll[RM9] mutant where cells all around the circumference adopt a lateral ectoderm (LE) fate. B Simulation of a tissue patch without a passive region. While active T1s deform the tension triangulation (top), the cell array (bottom) is blocked from elongating by the fixed DV “circumference” (implemented via slip walls). Instead, cell rearrangements are compensated by isogonal cell elongation (quantified in D) C CE (measured by the aspect ratio $a(t)$ of the active region) is strongly suppressed in absence of DV modulation of activity (“No DV”) and when active cells have vanishing shear modulus ($\mu_{a}=0$; see E and F). D Histograms of cell shape elongation measured by relative difference, $(S_{\text{DV,DV}}-S_{\text{AP,AP}})/(S_{\text{DV,DV}}+S_{\text{AP,AP}})$, of the AP-AP and DV-DV components of the shape tensor $S$. In the WT case (cf. Fig. 6), cells remain nearly isotropic while they become significantly oriented along the DV axis when DV modulation of activity or rigidity of active cells are abolished. E When the active cells have no shear rigidity ($\mu_{a}=0$), cell rearrangements are compensated by isogonal cell elongation (see D) without incurring an elastic energy build-up. Thus, almost no tissue-scale CE takes place. F Net amount of CE as a function of the shear modulus of active cells shows that active cells need to be stiffer than the surrounding passive tissue for active T1s to drive efficient CE.
@@
The total tissue extension found in the simulations that combine active and passive tissue regions is smaller than the extension of active tissue patches with free boundaries (compare Fig. 6D and Fig. 2C). This suggests that the passive tissue resists deformation. In the following, we further investigate the role of the spatial modulation of the cells’ mechanical properties along the DV axis. Figure 7B shows a simulation without DV modulation where all cells are active. Positive tension feedback drives active T1s everywhere, as is manifest in the deformation of the tension triangulation (Fig. 7B, right). However, because of the slip-wall boundary conditions, the tissue cannot contract along the DV axis so that T1s do not result in tissue convergence extension (see Movie 5; quantification in Fig. 7C). Instead, cell rearrangements are compensated by isogonal deformations resulting in elongated cell shapes (as quantified in Fig. 7E). We predict that this scenario will be realized in Toll[RM9] mutant embryos where all cells around the embryo’s circumference adopt a ventro-lateral fate [Irvine.Wieschaus1994], as illustrated in the cartoon in Fig. 7A. The stretching of cells leads to a buildup of elastic energy (see Fig. S10D). In Toll[RM9] mutant embryos, some of this elastic energy is released by the formation of folds (buckling) [Irvine.Wieschaus1994] which our 2D simulations cannot capture.
@@
Interestingly, the length dynamics of collapsing and emerging interfaces is not significantly affected by the lack of tissue-scale mechanical patterning (Fig. 7B), even though there is no tissue extension. This behavior has been observed in experiments and was interpreted as evidence for additional active mechanisms that drive interface elongation, e.g. such as medial myosin pulses [Collinet.etal2015]. However, interface extension in our model is a purely passive process, resulting from the temporal asymmetry of the intercalation process (i.e. the low level of active tension on the new interface), we conclude that no such mechanisms are necessary (see Fig. S10 for further details).
@@
The resistance of cells against shape deformations controls the isogonal modes and therefore is an important parameter controlling the interplay of active and passive tissue deformations. Fig. 7F shows a simulation where the shear modulus $\mu_{a}$ of active cells is set to zero. Active T1s can therefore be fully compensated by cell elongation through isogonal deformations without incurring an elastic energy cost. As a result, there is no net tissue deformation (see Movie 5). In other words, cell-shape rigidity is required to maintain rotund cell shapes (i.e. resist isogonal shear deformations) and thus translate active T1s into net tissue deformation. The tissue deformation by isogonal modes is determined by a balance of external forces and internal resistance of cells to shape changes. Here, the external forces acting on the active tissue result from the passive tissue’s resistance to deformation, which is in turn set by shear modulus $\mu_{\mathrm{p}}$. The ratio of the shear moduli in the active vs the passive region, $\mu_{\mathrm{a}}/\mu_{\mathrm{p}}$, determines how much the active region deforms (see Fig. 7G). Only when the cells in the active tissue are more rigid than those in the passive region ($\mu_{\mathrm{a}}/\mu_{\mathrm{p}}>1$), is it energetically favorable to isogonally deform the passive region rather than the active region. This predicts that GBE can be impaired by stiffening the dorsal tissue (amnioserosa), e.g. in dpp mutants.